OpenStudy (anonymous):

In trigonometry, cofunctions (sine and cosine, cosecant and secant, and tangent and cotangent) of __________________________ angles are congruent.

zepdrix (zepdrix):

Mmmmmm perpendicular angles? Is that what they're looking for maybe? \[\Large \sin(\theta+90^o)\quad=\quad \cos \theta\]